OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1). - -7/3 - 3/7 - 7/3 - -3/7

OpenStudy (anonymous):

To find slope you would do y2-y1/x2-x1. Once you do that you'll have you answer.

OpenStudy (anonymous):

i know, but when i did it, i got 7/-4

OpenStudy (anonymous):

1-(-6)/-3-(-6)= 1+6/-3+6= 7/3

OpenStudy (anonymous):

that's slope-intercept @Migitmack

OpenStudy (migitmack):

oh my bad

OpenStudy (anonymous):

Lol no prob. :)

OpenStudy (anonymous):

thank you.. :O ;D

OpenStudy (migitmack):

so whats the one 4 point slope i think it is y-y1=m(x-x1)

OpenStudy (migitmack):

oh he already said that-.-